my sister is 2 years older than i am. the sum of our ages is 36. how old are both of us.
Answer by T-Maree(17)

difference in age = d
d = 2 years
36 - d = 34
34 / 2 = 17
you are 17 and your sister is 19
17 + 19 = 36